OpenStudy (anonymous):

Solve for x: 4 - (x + 2) < -3(x + 4) can some one help me plz ? ^^

jigglypuff314 (jigglypuff314):

Hello and welcome to Openstudy! :) solve for x as you would for a normal equation 4 - (x + 2) < -3(x + 4) so first distribute the -3 4 - (x + 2) < -3x - 12 then distribute the - 4 - x - 2 < -3x - 12 add like terms. -x + 2 < -3x - 12 then you can add 3x to both sides +3x +3x 2x + 2 < -12 subtract 2 from both sides then divide both sides by 2

OpenStudy (anonymous):

thank you

jigglypuff314 (jigglypuff314):

glad I could help :)

OpenStudy (anonymous):

:D i m such a lost person hahaha

jigglypuff314 (jigglypuff314):

no worries :) many people are when it comes to math :P practice makes perfect ^.^

OpenStudy (primeralph):

It's not done though.

OpenStudy (anonymous):

indeed

jigglypuff314 (jigglypuff314):

I gave the last two steps that you need to do in order to get the answer. once you do them, tell me what you get ay?

OpenStudy (anonymous):

lost it lol

OpenStudy (primeralph):

I expected that.

jigglypuff314 (jigglypuff314):

2x + 2 < -12 subtract 2 from both sides then divide both sides by 2

OpenStudy (anonymous):

:( what?

jigglypuff314 (jigglypuff314):

alright :) let's start from 4 - (x + 2) < -3(x + 4) do you know how to distribute -3(x+4) = ?

OpenStudy (anonymous):

-3X+12

jigglypuff314 (jigglypuff314):

be careful! :) there's a negative attached to the 3 that carries through so -3(x+4) = (-3)(x) + (-3)(4) = -3x - 12 get it? :)

OpenStudy (anonymous):

X<-7

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

^^ CAN YOU HELP ME WITH ANOTHER ONE?

jigglypuff314 (jigglypuff314):

sure :)

OpenStudy (anonymous):

Solve for x: -3x + 3 < 6

jigglypuff314 (jigglypuff314):

subtract 3 from both sides then divide both sides by negative 3 ^ for this part don't for get to flip the sign (from < to >) because when you multiply or divide by a negative number then sign flips.

OpenStudy (anonymous):

x<-1

jigglypuff314 (jigglypuff314):

don't forget to flip the sign from < to >

OpenStudy (anonymous):

ahhhh ty

jigglypuff314 (jigglypuff314):

you're welcome :)

OpenStudy (anonymous):

Eduardo solved the following inequality and his work is shown below: -5(x + 4) + 21 ≥ -3 + 4(x - 8) -5x - 20 + 21 ≥ -3 + 4x - 32 -5x + 1 ≥ 4x - 35 -9x ≥ -36 x ≥ 1 What mistake did Eduardo make in solving the inequality? Select one: a. When dividing by -9, he did not change the ≥ to ≤. b. He subtracted 4x from both sides when he should have added 5x. c. He subtracted 1 from both sides when he should have added 36. d. He did not make a mistake.

jigglypuff314 (jigglypuff314):

it was the same mistake you had made in the previous problem :) when you multiply or divide by a negative number, the sign changes

OpenStudy (anonymous):

ohhh so its A

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

A bag of chips costs $3.79 including tax. Mr. Connor wants to purchase chips for his class and has a $15 budget. Write an inequality to solve for the number of bags of chips Mr. Connor can purchase with his budget. Select one: a. $15x ≥ $3.79 b. $15x ≤ $3.79 c. $3.79x ≥ $15 d. $3.79x ≤ $15

jigglypuff314 (jigglypuff314):

so it is $3.79 per bag of chips right? "per" means multiplying and if you replace "bag of chips" as "x" you would get "3.79x" alright? :)

jigglypuff314 (jigglypuff314):

then at most they can spend $15 so the value that they have has to be less than 15 "less than" = <

OpenStudy (anonymous):

$3.79x ≥ $15

OpenStudy (anonymous):

ohhhhh

OpenStudy (anonymous):

so its D

jigglypuff314 (jigglypuff314):

correct! :)

OpenStudy (anonymous):

your like awesome lol i m sorry if i asking to many questions

jigglypuff314 (jigglypuff314):

it's alright ;) I'm here because I like answering questions :P but there's this "Best Response" button over there --> It really encourages the helpers on this site to keep working for free xD

OpenStudy (anonymous):

Evelyn has $524.96 in her checking account. She must maintain a $500 balance to avoid a fee. She wrote a check for $32.50 today. Write and solve an inequality to solve for the least amount of money she needs to deposit to avoid a fee. Select one: a. 524.96 - 32.50 + x ≥ 500; x ≥ $7.54 b. 524.96 - 32.50 + x ≤ 500; x ≤ $7.54 c. 524.96 + 32.50 - x ≥ 500; x ≥ $57.46 d. 524.96 + 32.50 - x ≤ 500; x ≤ $57.46

OpenStudy (anonymous):

ohhh ok i will haha-clicks it like a boss-

jigglypuff314 (jigglypuff314):

"she must maintain a $500 balance" so her value has to be greater than 500 "greater than" = >

jigglypuff314 (jigglypuff314):

"wrote a check" would be subtracting money from her account so "- 32.50"

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

its A

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

whoooo hooo

OpenStudy (anonymous):

well thank you very much for helping these poor person here hehehe my name is Luxy by the way

jigglypuff314 (jigglypuff314):

Glad to I could help :) my name's Emily

OpenStudy (anonymous):

aww cute name

jigglypuff314 (jigglypuff314):

thanks :) I like yours too :D

OpenStudy (anonymous):

:D

jigglypuff314 (jigglypuff314):

if you ever need help on anymore questions you can tag them by putting a "@" in front of their name :) like @NatashaMoon

OpenStudy (anonymous):

do u have facebook?

jigglypuff314 (jigglypuff314):

yes I do, but I would prefer to not share it :P if you want to contact me outside of this site my email is jigglypuff314@gmail.com

OpenStudy (anonymous):

i m not raper lolzzz i just like new friends but ok ^-^ i ma put it down

jigglypuff314 (jigglypuff314):

if you like making new friends and socializing, there's a chat box in the bottom right corner of the screen :)

OpenStudy (anonymous):

i know i saw it O_O they sure talk fast mhm

jigglypuff314 (jigglypuff314):

indeed lol

OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) + 6.

jigglypuff314 (jigglypuff314):

3(x + 1) = -2(x - 1) + 6 distribute the 3 3x + 3 = -2(x - 1) + 6 distribute the negative 2 3x + 3 = -2x + 1 + 6 add like terms 3x + 3 = -2x + 7 add 2x to both sides subtract 3 from both sides divide both sides by 5 it should end up as a decimal or fraction :)

OpenStudy (anonymous):

thanx :D

OpenStudy (anonymous):

A store offers a 15% discount on all items in the store during a sale. All store employees get an additional 10% employee discount. In the equation below, C represents the cost of an item for an employee including tax, and x represents the original price of the item. C = 0.765x + 0.06(0.765x) Chris is an employee at the store. He plans to buy a coat during the sale. The price of the coat before the discount is $100. How much will Chris pay for the coat during the sale?

jigglypuff314 (jigglypuff314):

so you would plug in 100 for x 0.765(100) + (0.06)(0.765)(100) =

OpenStudy (anonymous):

so u add 0.765 to 100 ?

jigglypuff314 (jigglypuff314):

when parenthesizes are next to each other like that, it means multiplication.

OpenStudy (anonymous):

so i multipli (100)+(0.06)?

OpenStudy (anonymous):

then do the same for the other?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

so for the first one i got 6 fotr the other 76.5

jigglypuff314 (jigglypuff314):

hmm... close. :) 0.765 times(100) = 76.5 (0.06)times(0.765)times(100) = 4.59 then add the two numbers :)

OpenStudy (anonymous):

81.09?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

so that is the answer?

jigglypuff314 (jigglypuff314):

yup :)

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

ty

jigglypuff314 (jigglypuff314):

I've got to go now... Have a Nice day! and I'll see ya around :)

OpenStudy (anonymous):

thank you for helping ^^